OpenStudy (anonymous):

what is the value of log5 25? 2 5 0.699 1.398

OpenStudy (danjs):

the base to what power is 25

OpenStudy (danjs):

\[\huge \log_{b} x = y~~~or~~~b^y=x\]

OpenStudy (anonymous):

it says log5 and the 5 is little, I just don't know what symbol to use to show it

OpenStudy (anonymous):

Well first we use the power of love to mathematically solve.

OpenStudy (anonymous):

Then you will find the answer deep within your heart.

OpenStudy (anonymous):

Got that? Write that down.

OpenStudy (anonymous):

what number do you raise 5 to the power of to get 25?

OpenStudy (anonymous):

love

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

Charolette I really don't think that's funny. In genuinely trying to get help so get off my problem

OpenStudy (danjs):

\[\huge \log_{5} \] when you see this 'log', although it could be l*o*g, think of it as a sentence that reads, "What power do i have to raise this small number to in order to get" Here is says,"what power do i have to raise 5 to, in order to get"

OpenStudy (danjs):

5^(power) = 25 5*5 = 25, you need 2 of em , 5^2=25

OpenStudy (anonymous):

Okay, so the answers 5? I tried that and I got a huge number

OpenStudy (danjs):

\[\huge \log_{5}25 = 2 \]

OpenStudy (danjs):

5 to the second power is 25

OpenStudy (anonymous):

Ohhh, so it's 2! Okay thank you! I thought that in the beginning but wanted to double check :) thank you again!

OpenStudy (danjs):

What power do you have to raise 5 to, to get 25?", answer, 2

OpenStudy (danjs):

welcome